OpenStudy (anonymous):

I need Help with gas calculations, I think its stoichiometry. I Know how to do it when they give me a question like this "How many grams of water vapor (H2O) are in a 10.2 liter sample at 0.98 atmospheres and 26°C?" But questions such as this "How many liters of water vapor can be produced if 13.3 liters of methane gas (CH4) are combusted, if all measurements are taken at the same temperature and pressure? CH4 (g) + 2 O2 (g) CO2 (g) + 2 H2O (g)" Confuse me, even though they don't look that much different. Does anyone know how to do this stuff an can maybe explain or give me a link?

OpenStudy (aaronq):

For the latter question, all you need to do it work with the stoichiometric coefficients. Consider, a general reaction where a, b, c and d are the coefficients and A, B, C and D are the species. aA + bB -> cC + dD oftentimes you can reason it out, but building a mathematical relationship will not ever fail. \[\frac{ n _{A} }{ a }=\frac{ n _{B}}{ b }=\frac{ n _{C} }{ c }=\frac{ n _{D} }{ d }\] so you can use just the part you're interested in, say you know the moles of B equal to 2, and you're wondering what the moles of D will be (assuming the reaction went to completion): \[\frac{ n _{B} }{ b }=\frac{ n _{D} }{ d }\] substitute your values, then isolate (algebraically) what you want (your unknown): \[\frac{ 2 }{ b }=\frac{n _{D} }{ d } \rightarrow n _{D}=\frac{ 2d }{ b }\]

OpenStudy (anonymous):

...Is there a version of this for dumb people?

OpenStudy (aaronq):

lol i thought what i wrote was pretty straight forward... il give you an example maybe that will clarify things. say you have reaction: CH4 + 2O2 -> CO2 + 2H2O they ask you to "find how many moles (or grams all you need to do is convert) of H2O are produced when you burn 4 moles of CH4?" all you do is build a ratio: \[\frac{ n _{CH _{4}} }{ 1}=\frac{ n _{H _{2}O} }{ 2 }\] *notice their respective coefficients are underneath their moles: \[\frac{ 4 }{ 1 }=\frac{ n _{H _{2}O} }{ 2 }\] then isolate the moles of water: \[n _{H _{2}O}=\frac{ 2(4) }{ 1 } = 8\] and so you produced 8 moles of water. is that clearer?

OpenStudy (anonymous):

Yes, Thank you so much. The whole analogy thing made it more confusing. Thank you for taking your time and writing out everything

OpenStudy (aaronq):

no problem, dude. At first, I thought maybe giving you a general abstract example would help but i guess not.

OpenStudy (anonymous):

ya it helps now that I see the example, but before the example it was to much to try and put the two and two together. What can I say Im just bad at chem xD

OpenStudy (aaronq):

haha it's cuz it's mathy. anyway, good stuff you understand it, maybe you can explain it one day to the many people that ask this same question over and over and over.. lol